# Learning non-Gaussian spatial distributions via Bayesian transport maps with parametric shrinkage

### v1

@@ Abstract
Many applications, including climate-model analysis and stochastic weather generators, require learning or emulating the distribution of a high-dimensional and non-Gaussian spatial field based on relatively few training samples. To address this challenge, a recently proposed Bayesian transport map (BTM) approach consists of a triangular transport map with nonparametric Gaussian-process (GP) components, which is trained to transform the distribution of interest distribution to a Gaussian reference distribution. To improve the performance of this existing BTM, we propose to shrink the map components toward a “base” parametric Gaussian family combined with a Vecchia approximation for scalability. The resulting ShrinkTM approach is more accurate than the existing BTM, especially for small numbers of training samples. It can even outperform the “base” family when trained on a single sample of the spatial field. We demonstrate the advantage of ShrinkTM though numerical experiments on simulated data and on climate-model output.
@@ 1 Introduction
We propose a novel extension of the BTM approach of Katzfuss and Schäfer, (2024) that is specifically tailored for learning the distributions of large spatial fields when very limited training samples (often, only one) are available. This extension builds on Kidd and Katzfuss, (2022), who proposed Bayesian nonparametric inference on the inverse Cholesky factors of the GP covariance matrices and empirically studied their parametric regularization. The major contribution lies in defining a new set of prior distributions for the map components that first centers the mean and variance of each of the components toward the conditional means and variances of a parametric GP (e.g., with a Matérn covariance). Next, it introduces some regularization factors in the prior distributions, which can be optimized to actively learn the amount of shrinkage towards the GP family. We approximate the conditional means and variances by leveraging Vecchia approximation Vecchia, (1988). The resulting regularized BTM approach, which we call ShrinkTM, is fast and produces accurate inference for large spatial non-Gaussian fields even when trained with only $n=1$ sample. Consequently, it expands the application of the BTM approach introduced by Katzfuss and Schäfer, (2024) in many realistic applications.
The remainder of the paper is organized as follows. In Section 2, we review the Bayesian spatial transport map of Katzfuss and Schäfer, (2024). In 3, we introduce our regularization for the BTM approach, including the use of the Vecchia approximation for computing conditional means and variances of the base GP family. In Section 4, we provide numerical comparisons of our methods to the existing BTM approach using simulated data and climate-model output. We conclude in Section 5. Proofs are provided in Appendix A.
@@ 2.3 Maximin ordering for large spatial fields
When the variables $y_{1},\ldots,y_{N}$ are not associated with spatial locations or when Euclidean distance between the locations is not meaningful (e.g., nonstationary, multivariate, spatio-temporal, or functional data), the maximin and neighbor ordering can be carried out based on other distance metrics, such as $(1-|\text{correlation}|)^{1/2}$ based on some guess or estimate of the correlation between variables (Kang and Katzfuss, , 2023; Kidd and Katzfuss, , 2022).
@@ 2.4 Computation and inference for large spatial fields
$q_{i,c_{i}(k)}=\begin{cases}\exp(\theta_{q}k),&k\leq m^{\prime},\\ 0,&k>m^{\prime}.\end{cases}$ (7)
The sparsity parameter $m^{\prime}=\max\{k:\exp(\theta_{q}k)\geq\varepsilon\}$ is determined by the data through the hyperparameter $\theta_{q}$, where $\varepsilon$ is a small fixed threshold (e.g., $\varepsilon=0.01$). The parametrization of $q_{i,c_{i}(k)}$ has been primarily considered by Katzfuss and Schäfer, (2024) and is motivated by exponential rates of screening for Gaussian processes derived from elliptic boundary-value problems (Schäfer et al., 2021b, ; Schäfer et al., 2021a, ; Kang et al., , 2024). Consequently, $f_{i}(\mathbf{y}_{1:i-1})=f_{i}(\mathbf{y}_{g_{m^{\prime}}(i)})$ and its covariance kernel can be expressed as
@@ 3.2 Nonlinear prior distributions
Similar to (5), we model each function $f_{i}$ as a GP conditional on $d_{i}^{2}$:
@@
where $\gamma$, $h_{i}$ and $\rho$ follow similar parameterization as in (8). To obtain the sparsity parameter $m^{\prime}$ in (7), we used $\varepsilon=0.01$ for our numerical examples, which produced highly accurate inference and usually resulted in $m^{\prime}<10$. In contrast to (8), we introduce a shrinkage parameter $\sigma^{2}_{0}$ in (13) that decides the level of shrinkage towards a Gaussian process regression.
@@ 4.1 Simulation experiments
We consider both the existing BTM approach of Katzfuss and Schäfer, (2024) (which we name SimpleTM) and our new version (i.e., ShrinkTM). For each of the two simulation settings, we randomly sample $n$ training samples and train both SimpleTM and ShrinkTM. We vary the value of $n$ and repeat this experiment 10 times for each value of $n$. We perform the training of both SimpleTM and ShrinkTM in Python using the stochastic-gradient-descent-based Adam optimization algorithm from the PyTorch library. We also compare to a MatCov approach that refers to a Gaussian process with an isotropic Matérn covariance (which is also used as the base covariance in ShrinkTM), with the three hyperparameters inferred via maximum likelihood estimation.
@@
Figures 2 and 3 show samples generated from SimpleTM and ShrinkTM for different training sizes for LR900 and NR900, respectively. We can see that even for $n=1$ training samples, ShrinkTM excels in capturing the long and short range dependencies due to better regularization through the conditional means and variances, whereas SimpleTM struggles. We can furthermore see in Figure 5(a) that for LR900, the shrinkage factor $c_{d}$ lies between 0.1 and 0.25 for $n=1$ with the average value of $c_{d}$ around 0.15, indicating the efficiency of ShrinkTM. This average gradually drops to 0 as the sample size increases. As $c_{d}<1$ means $sd(d_{i})<E(d_{i})$, this suggests a moderately strong shrinkage of the variance components $d_{i}$ toward the corresponding values $\tau_{i}(\hat{\theta}_{p})$’s (as discussed in Section 3.2); this shrinkage increases with $n$, as $c_{d}$ decreases more. On the other hand, since NR900 has a modified dependence structure relative to an exponential covariance, we can see that $c_{d}$ varies for different sample sizes and automatically decides the amount of shrinkage from the training data.
@@
To further assess the accuracy of these two methods, we generate test samples from the two simulation scenarios, and we compare average logarithmic score (higher is better) of SimpleTM, ShrinkTM, and MatCov in Figure 5. For LR900, the log-score of ShrinkTM is much higher than that of SimpleTM for very small numbers of training samples $n$ (as seen in Figure 5(a)), and despite its flexible nonparametric structure, it matches the accuracy of the correctly specified parametric “base” MatCov (which only needs to estimate three parameters) when trained on only $n=2$ training samples. Even when data is being generated from NR900, ShrinkTM consistently outperforms SimpleTM (as seen in Figure 5(b)). Training time of ShrinkTM was also comparable with that of SimpleTM: On the uniform grid of $N=900$ points, ShrinkTM took 49 seconds and SimpleTM took 7 seconds to train on $n=1$ replicate, but ShrinkTM and SimpleTM took 14.5 and 16 minutes, respectively, when training on $n=200$ samples. This is in line with the time complexities of $\mathcal{O}(Nm^{3})$ for the Vecchia approximation (only in ShrinkTM) and $\mathcal{O}(N(n^{3}+m^{\prime}n^{2}))$ for computing and decomposing the $\mathbf{G}_{i}$ (for both ShrinkTM and SimpleTM), where the latter increases and starts to dominate for increasing $n$.
@@ 4.2 Climate data application
We consider log-transformed total precipitation rate (in m/s) on a roughly $1^{\circ}$ longitude-latitude global grid of size N = 37 × 74 = 2738 in the middle of the Northern summer(July 1) in 98 consecutive years, starting in the year 402, from the Community Earth System Model (CESM) Large Ensemble Project (Kay et al., (2015)). This dataset was also analyzed in Katzfuss and Schäfer, (2024). CESM belongs to a broader class of climate models, which are large sets of computer code describing the behavior of the Earth system (e.g., the atmosphere) via systems of differential equations. Enormous computational power is required to produce even a single ensemble of these models on fine latitude-longitude grids from these computer codes. Due to the massive cost of obtaining a single sample from these climate models, it is important to build a stochastic generative model based on few training samples from the climate model, in order to produce relevant summaries and even draw more samples at much cheaper cost.
@@
We considered a subsection of the western hemisphere (39.1N to 29.6S and 110W to 65W) containing parts of land including North, Central and South America, and a subsection of Atlantic and Pacific ocean. We obtained precipitation anomalies by standardizing the data at each grid location to mean zero and variance 1 (as shown in Figure 6(a)), and used them as training ensembles.
Figure 6 also displays a set of samples obtained from SimpleTM, ShrinkTM, and MatCov for several training ensemble sizes. It is evident that with a low number of training samples (i.e., $n=1$), ShrinkTM is able to capture non-Gaussian features and long-range characteristics, while SimpleTM requires 10 or more training samples to exhibit such features. Training of ShrinkTM takes 39 seconds in comparison to 3 seconds taken by SimpleTM for $n=1$. The training time increases to 9, 18, 82 seconds for SimpleTM, and 46, 53, 123 seconds for ShrinkTM when training on $n=10,20,50$ samples, respectively. Figure 7(a) shows that for small $n$, the shrinkage factor $c_{d}$ has an extremely high value, suggesting non-Gaussianity in the samples. The level of non-Gaussianity is better learned with increase in training ensembles.
We further used the idea of conditional simulation to sample precipitation anomalies when a sample is partially observed. To do this, we consider the precipitation anomalies in the top-left panel of Figure 6 (which were not part of the training data) and assume that we have observed only the first $100$ ordered locations of this test field. Figure 8 displays the conditional samples for different methods for varying training ensembles conditional on this partially observed field. It can be seen (Figure 8(b)) that with only a few number of training ensembles, ShrinkTM is able to capture most of the long range dependencies that are seen in Figure 6 (topleft panel) as well. SimpleTM needs around 30 samples to capture these dependencies, while MatCov is not able to capture them at all.
We further compare the performance of ShrinkTM and SimpleTM using average log-score. Figure 7(c) shows the average log-score for threse three methods with varying number of training ensembles. ShrinkTM outperforms both SimpleTM and MatCov for all training-sample sizes, including providing higher accuracy than the “base” MatCov approach even for a single training sample.
@@ 5 Conclusion
We proposed an improved version of the BTM (Katzfuss and Schäfer, (2024)) by centering the mean and variance functions of its map components around the conditional means and variances obtained from a Gaussian process (GP) with a parametric covariance, whose parameters are estimated from the training data. We also employ a Vecchia approximation in the calculation of conditional means and variances of the GP. This new BTM version, which we call ShrinkTM, preserves the scalability and flexibility of BTM. A notable advantage of ShrinkTM is that it can cheaply produce realistic samples even when the transport map is trained with only a single training ensemble, and it is thus especially useful when few or even only a single training sample is available. ShrinkTM is nonparametric and largely automated, in that it does not require specifying the form or extent of non-Gaussianity or nonstationarity. We have demonstrated these crucial advantages using simulated datasets and log-precipitation datasets obtained from the CESM climate model. A Python implementation of our methods, along with code to reproduce all results, is available at https://github.com/katzfuss-group/batram/tree/ShrinkTM. Our work can be further extended to multivariate transport maps (e.g., Wiemann and Katzfuss, (2023)). Moreover, we can vary the hyperparameters of the Gaussian process over a covariate space and thus produce a covariate-dependent shrinkage of the heteroscedastic transport map Drennan et al., prep , which can be applicable to a wider range of applications. We plan to tackle these problems in future work.

### v2

@@ Abstract
Many applications, including climate-model analysis and stochastic weather generators, require learning or emulating the distribution of a high-dimensional and non-Gaussian spatial field based on relatively few training samples. To address this challenge, a recently proposed Bayesian transport map (BTM) approach consists of a triangular transport map with nonparametric Gaussian-process (GP) components, which is trained to transform the distribution of interest to a Gaussian reference distribution. To improve the performance of this existing BTM, we propose to shrink the map components toward a “base” parametric Gaussian family combined with a Vecchia approximation for scalability. The resulting ShrinkTM approach is more accurate than the existing BTM, especially for small numbers of training samples. It can even outperform the “base” family when trained on a single sample of the spatial field. We demonstrate the advantage of ShrinkTM through numerical experiments on simulated data and on climate-model output.
@@ 1 Introduction
We propose a novel extension of the BTM approach of Katzfuss and Schäfer, (2024) that is specifically tailored for learning the distributions of large spatial fields when very limited training samples (often, only one) are available. This extension builds on Kidd and Katzfuss, (2022), who proposed Bayesian nonparametric inference on the inverse Cholesky factors of the GP covariance matrices and empirically studied their parametric regularization. The major contribution lies in defining a new set of prior distributions for the map components that first centers the mean and variance of each of the components toward the conditional means and variances of a parametric GP (e.g., with a Matérn covariance). Next, it introduces some regularization factors in the prior distributions, which can be optimized to actively learn the amount of shrinkage towards the GP family. We approximate the conditional means and variances by leveraging Vecchia approximation (Vecchia, , 1988). The resulting regularized BTM approach, which we call ShrinkTM, is fast and produces accurate inference for large spatial non-Gaussian fields even when trained with only $n=1$ sample. Consequently, it expands the application of the BTM approach introduced by Katzfuss and Schäfer, (2024) in many realistic applications.
The remainder of the paper is organized as follows. In Section 2, we review the Bayesian spatial transport map of Katzfuss and Schäfer, (2024). In Section 3, we introduce our regularization for the BTM approach, including the use of the Vecchia approximation for computing conditional means and variances of the base GP family. In Section 4, we provide numerical comparisons of our methods to the existing BTM approach using simulated data and climate-model output. We conclude in Section 5. Proofs are provided in Appendix A.
@@ 2.3 Maximin ordering for large spatial fields
When the variables $y_{1},\ldots,y_{N}$ are not associated with spatial locations or when Euclidean distance between the locations is not meaningful (e.g., nonstationary, multivariate, spatio-temporal, or functional data), the maximin and neighbor ordering can be carried out based on other distance metrics, such as $(1-|\text{correlation}|)^{1/2}$ based on some guess or estimate of the correlation between variables (Kidd and Katzfuss, , 2022; Kang and Katzfuss, , 2023).
@@ 2.4 Computation and inference for large spatial fields
$q_{i,c_{i}(k)}=\begin{cases}\exp(-e^{\theta_{q}}k),&k\leq m^{\prime},\\ 0,&k>m^{\prime}.\end{cases}$ (7)
The sparsity parameter $m^{\prime}=\max\{k:\exp(-e^{\theta_{q}}k)\geq\varepsilon\}$ is determined by the data through the hyperparameter $\theta_{q}$, where $\varepsilon$ is a small fixed threshold (e.g., $\varepsilon=0.01$). The parametrization of $q_{i,c_{i}(k)}$ has been primarily considered by Katzfuss and Schäfer, (2024) and is motivated by exponential rates of screening for Gaussian processes derived from elliptic boundary-value problems (Schäfer et al., 2021b, ; Schäfer et al., 2021a, ; Kang et al., , 2024). Consequently, $f_{i}(\mathbf{y}_{1:i-1})=f_{i}(\mathbf{y}_{g_{m^{\prime}}(i)})$ and its covariance kernel can be expressed as
@@ 3.2 Nonlinear prior distributions
Conditional on the variance component $d_{i}^{2}$, we model each function $f_{i}$ using a GP:
@@
where $\gamma$, $h_{i}$ and $\rho$ follow similar parameterization as in (8). To obtain the sparsity parameter $m^{\prime}$ in (7), we used $\varepsilon=0.01$ for our numerical examples, which produced highly accurate inference and usually resulted in $m^{\prime}<10$. In contrast to (8), we introduce a shrinkage parameter $\sigma^{2}_{0}$ in (13) that determines the level of shrinkage towards a linear regression.
@@
Figure 2: For two locations at distance 0.3 in the Gaussian simulation example in Figure 3, the true bivariate Gaussian distribution (contour lines), along with 5,000 samples each from the fitted existing Bayesian transport map (SimpleTM) outlined in Section 2 (red) and our proposed ShrinkTM (blue), both trained on a single ($n=1$) sample.
Equations (10)–(13) imply a transport-map model that we call ShrinkTM, which can accurately learn the joint distribution even when trained with small number of samples. For example, Figure 2 shows that if we train both ShrinkTM and the existing transport-map approach (SimpleTM) outlined in Section 2 on single simulated sample, ShrinkTM’s parametric shrinkage allows it to capture the true distribution among two arbitrarily selected locations much more accurately than SimpleTM, which produces overly heavy tails.
@@ 4.1 Simulation experiments
We consider both the existing BTM approach of Katzfuss and Schäfer, (2024) (which we name SimpleTM) and our new version (i.e., ShrinkTM). For each of the two simulation settings, we randomly sample $n$ training samples and train both SimpleTM and ShrinkTM. We vary the value of $n$ and repeat this experiment 10 times for each value of $n$. For all the training tasks, we use default initial values $\bm{\theta}=(\bm{\theta}_{p},c_{d},\theta_{\sigma,1},\theta_{\sigma,2},\theta_{q})=(2.0,0.0,0.0,0.0,-1.0)$. We perform the training of both SimpleTM and ShrinkTM in Python using the stochastic-gradient-descent-based Adam optimization algorithm from the PyTorch 2.5.1 library. We set the initial learning rate of the Adam optimizer to be 0.01 and adjust it using the cosine annealing rule in the PyTorch library. The remaining tuning parameters of the optimizer are fixed at their default values. We also compare to a MatCov approach that refers to a Gaussian process with an isotropic Matérn covariance (which is also used as the base covariance in ShrinkTM), with the three hyperparameters inferred via maximum likelihood estimation.
@@
Figures 3 and 4 show samples generated from SimpleTM and ShrinkTM for different training sizes for LR900 and NR900, respectively. We can see that even for $n=1$ training samples, ShrinkTM excels in capturing the long- and short-range dependencies due to better regularization through the conditional means and variances, whereas SimpleTM struggles. We can furthermore see in Figure 5(a) that for LR900, the shrinkage factor $c_{d}$ lies between 0.1 and 0.25 for $n=1$ with the average value of $c_{d}$ around 0.15, indicating the efficiency of ShrinkTM. As $c_{d}<1$ means $sd(d_{i})<E(d_{i})$, this suggests a moderately strong shrinkage of the variance components $d_{i}$ toward the corresponding values $\tau_{i}(\hat{\theta}_{p})$’s (as discussed in Section 3.2). The average of $c_{d}$ decreases for large training sample sizes, indicating a stronger shrinkage. On the other hand, since NR900 has a modified dependence structure relative to an exponential covariance, we can see that $c_{d}$ varies for different sample sizes and automatically decides the amount of shrinkage from the training data.
@@
To further assess the accuracy of these two methods, we generate test samples from the two simulation scenarios, and we compare average logarithmic score (higher is better) of SimpleTM, ShrinkTM, and MatCov in Figure 6. The log-score (e.g., Gneiting2014) approximates the negative Kullback-Leibler divergence between the true and approximated distribution up to an additive constant. For LR900, the log-score of ShrinkTM is much higher than that of SimpleTM for very small numbers of training samples $n$ (as seen in Figure 6(a)), and despite its flexible nonparametric structure, it matches the accuracy of the correctly specified parametric “base” MatCov (which only needs to estimate three parameters) when trained on only $n=2$ training samples. Even when data is being generated from NR900, ShrinkTM consistently outperforms SimpleTM (as seen in Figure 6(b)). Average training time of ShrinkTM was also comparable with that of SimpleTM: On the uniform grid of $N=900$ points, on average, ShrinkTM took 49 seconds and SimpleTM took 7 seconds to train on $n=1$ replicate, but ShrinkTM and SimpleTM took 15 and 13.5 minutes, respectively, when training on $n=200$ samples. This is in line with the time complexities of $\mathcal{O}(Nm^{3})$ for the Vecchia approximation (only in ShrinkTM) and $\mathcal{O}(N(n^{3}+m^{\prime}n^{2}))$ for computing and decomposing the $\mathbf{G}_{i}$ (for both ShrinkTM and SimpleTM), where the latter increases and starts to dominate for increasing $n$.
4.2 Climate data application
@@
We consider log-transformed total precipitation rate (in m/s) on a roughly $1^{\circ}$ longitude-latitude global grid of size N = 37 × 74 = 2738 in the middle of the Northern summer(July 1) in 98 consecutive years, starting in the year 402, from the Community Earth System Model (CESM) Large Ensemble Project (Kay et al., , 2015). This dataset was also analyzed in Katzfuss and Schäfer, (2024). CESM belongs to a broader class of climate models, which are large sets of computer code describing the behavior of the Earth system (e.g., the atmosphere) via systems of differential equations. Enormous computational power is required to produce even a single ensemble of these models on fine latitude-longitude grids from these computer codes. Due to the massive cost of obtaining a single sample from these climate models, it is important to build a stochastic generative model based on few training samples from the climate model, in order to produce relevant summaries and even draw more samples at much cheaper cost.
@@
We consider a subsection of the western hemisphere (39.1N to 29.6S and 110W to 65W) containing parts of land including North, Central and South America, and a subsection of Atlantic and Pacific ocean. We obtain precipitation anomalies by standardizing the data at each grid location to mean zero and variance 1 (as shown in Figure 7(a)), and used them as training ensembles.
Figure 10: In the partially observed setting of Figure 9, root mean square error at held-out locations for conditional samples from three generative models: SimpleTM, ShrinkTM, and GP with Matèrn kernel.
Figure 7 displays a set of samples obtained from SimpleTM, ShrinkTM, and MatCov for several training ensemble sizes. It is evident that with a low number of training samples (i.e., $n=1$), ShrinkTM is able to capture non-Gaussian features and long-range characteristics, while SimpleTM requires 10 or more training samples to exhibit such features. When trained on $n=1,5,10,30,50$ training samples, ShrinkTM selects $m^{\prime}=3,3,4,16,20$ nearest neighbors to learn non-Gaussian distribution, respectively. On average, training of ShrinkTM takes 39 seconds, in comparison to 3 seconds taken by SimpleTM for $n=1$. The average training time increases to 9, 18, 82 seconds for SimpleTM, and 46, 53, 123 seconds for ShrinkTM when training on $n=10,20,50$ samples, respectively. Figure 8(a) shows that for small $n$, the shrinkage factor $c_{d}$ has an extremely high value, suggesting non-Gaussianity in the samples. The level of non-Gaussianity is better learned with increase in training-ensemble size.
We quantitatively compare the performance of ShrinkTM and SimpleTM using average log-score. Figure 8(c) shows the average log-score for these three methods with varying number of training ensembles. ShrinkTM outperforms both SimpleTM and MatCov for all training-sample sizes, including providing higher accuracy than the “base” MatCov approach even for a single training sample.
We consider conditional simulation to sample precipitation anomalies when a sample is partially observed. To do this, we consider the precipitation anomalies in the top-left panel of Figure 7 (which were not part of the training data) and assume that we have observed only the first $100$ ordered locations of this test field. Figure 9 displays the conditional samples for different methods for varying training ensembles conditional on this partially observed field. It can be seen (Figure 9(b)) that with only a few number of training ensembles, ShrinkTM is able to capture most of the long range dependencies that are seen in Figure 7 (topleft panel) as well. SimpleTM needs around 30 samples to capture these dependencies, while MatCov is not able to capture them at all. Figure 10 shows root mean square error (RMSE) of predicting the held-out locations when repeating this experiment 10 times.
@@ 5 Conclusion
We proposed an improved version of the BTM (Katzfuss and Schäfer, , 2024) by centering the mean and variance functions of its map components around the conditional means and variances obtained from a Gaussian process (GP) with a parametric covariance, whose parameters are estimated from the training data. We also employ a Vecchia approximation in the calculation of conditional means and variances of the GP. This new BTM version, which we call ShrinkTM, preserves the scalability and flexibility of BTM. A notable advantage of ShrinkTM is that it can cheaply produce realistic samples even when the transport map is trained with only a single training ensemble, and it is thus especially useful when few or even only a single training sample is available. ShrinkTM is nonparametric and largely automated, in that it does not require specifying the form or extent of non-Gaussianity or nonstationarity. We have demonstrated these crucial advantages using simulated datasets and log-precipitation datasets obtained from the CESM climate model. A Python implementation of our methods, along with code to reproduce all results, is available at https://github.com/katzfuss-group/batram/tree/ShrinkTM. Our work can be further extended to multivariate transport maps (e.g., Wiemann and Katzfuss, (2023)). Moreover, we can vary the hyperparameters of the Gaussian process over a covariate space and thus produce a covariate-dependent shrinkage of the heteroscedastic transport map (Drennan et al., prep, ), which can be applicable to a wider range of applications. We plan to tackle these problems in future work.